# Event Data Association via Robust Model Fitting for Event-based Object Tracking

### v2

@@ Abstract
Event-based approaches, which are based on bio-inspired asynchronous event cameras, have achieved promising performance on various computer vision tasks. However, the study of the fundamental event data association problem is still in its infancy. In this paper, we propose a novel Event Data Association (called EDA) approach to explicitly address the event association and fusion problem. The proposed EDA seeks for event trajectories that best fit the event data, in order to perform unifying data association and information fusion. In EDA, we first asynchronously fuse the event data based on its information entropy. Then, we introduce a deterministic model hypothesis generation strategy, which effectively generates model hypotheses from the fused events, to represent the corresponding event trajectories. After that, we present a two-stage weighting algorithm, which robustly weighs and selects true models from the generated model hypotheses, through multi-structural geometric model fitting. Meanwhile, we also propose an adaptive model selection strategy to automatically determine the number of the true models. Finally, we use the selected true models to associate and fuse the event data, without being affected by sensor noise and irrelevant structures. We evaluate the performance of the proposed EDA on the object tracking task. The experimental results show the effectiveness of EDA under challenging scenarios, such as high speed, motion blur, and high dynamic range conditions.
@@
Event camera, Data association, Object tracking, Model fitting, Event fusion
journal: Journal of Information Fusion
@@ 1 Introduction
As one of the most delicate neural systems, biological retinas can precisely and efficiently capture object motions olveczky2003segregation . Inspired by the biological retina, asynchronous event cameras, such as DVS lichtsteiner2008128 , DAVIS brandli2014240 and ATIS posch2011qvga , have been proposed to mimic it. These bio-inspired silicon retina vision sensors have a very high dynamic range (higher than 100 dB) and an ultra-low latency (less than 1 ms). In particular, each pixel of the event camera can independently and asynchronously emit binary (i.e., On or Off) retinal events in responding to pixel intensity changes in the current environment. Pixel intensity changes are usually triggered by object and camera motions, under relatively smooth external illumination variations. Thus, the retinal events contain clear motion information. Moreover, thanks to the high dynamic range and low latency, event cameras are not affected by those challenges that inherently affect conventional frame-based cameras, such as motion blur and pixel saturation. As a result, event cameras are well suited for motion-related computer vision tasks.
@@
Event-based methods have achieved promising performance on various tasks gallego2022event . However, the study of the fundamental event data association problem is still challenging and in its infancy. Unlike a traditional camera, an event camera only sparsely emits binary (i.e., On and Off) retinal events at the edges within a scene, in responding to camera and object motions. Therefore, without the pixel intensity information, it is difficult to calculate event correspondences (i.e., event trajectories) in the temporal domain for event-based data association and fusion. Since data association is the key to various fundamental computer vision tasks (such as object tracking), the problem of event-based data association needs further investigation.
As the pioneering work on event-based data association, gallego2018unifying proposes a contrast maximization framework to implicitly handle the data association problem based on the corresponding motion parameters. In that framework, the initial motion parameters for each event are empirically set to a range of predefined values for optimization. Since the predefined motion range may not cover all possible event motions, the final result may not be the optimal estimate.
However, gallego2018unifying and other event-based data association methods zhu2017event ; gallego2019focus ; peng2022globally show that the events triggered by the same edge in the scene can be associated with each other using an event trajectory. Furthermore, they also show that the event trajectories, triggered at different edges of the scene, are nearly 3D curves in the spatio-temporal domain (see an example in Fig. 1). Based on these observed facts, we explicitly formulate the event-based data association problem as a 3D event trajectory estimation problem in the spatio-temporal domain. Then, a novel event-based data association (called EDA) approach is proposed to perform explicit and accurate event-based data association and fusion. In the proposed EDA, the event trajectory estimation problem is effectively solved through the theory of robust multi-structural model fitting. To the best of our knowledge, it is the first time to introduce robust model fitting for event-based data association. The estimated multiple structures correspond to the event trajectories triggered by different motions, and they are used to associate the event data. To evaluate the effectiveness of the proposed EDA, we use EDA to deal with the problem of event-based data association in object tracking scenarios. Overall, this paper makes the following contributions:
  • 1.
    We introduce a deterministic strategy to effectively generate model hypotheses in the spatio-temporal domain, spanned by retinal events, which are considered as event trajectory candidates for event data association.
  • 2.
@@ 2 Related Work
As mentioned previously, event cameras have shown their great potential on various computer vision tasks gallego2022event . Different from conventional cameras that are synchronized to the global camera shutter, these bio-inspired vision sensors work in an asynchronous way with ultra low microsecond latencies in responding to object motions, which is inspired by the biological retina. Thanks to the asynchronous nature, the high spatio-temporal resolution, and the superior HDR properties of event cameras, various event-based methods have been proposed for many computer vision tasks, especially for object tracking.
In the field of object tracking, conventional object tracking methods can be roughly categorized into three groups: correlation tracking methods, deep tracking methods, and hybrid tracking methods. For the correlation tracking methods, Liang2021robust and wang2021transformer exploit spatio-temporal information for discrimination tracking. xu2020accelerated uses accelerated correlation filters to track objects. liu2023visual fuses visual cognition flows for correlation tracking. In the deep tracking methodsmarvasti2022deep , cui2022joint leverages a classification and regression Siamese framework for deep tracking. chen2022siamban adopts a Siamese box adaptive network for target-aware tracking. mayer2022transforming and cui2022mixformer respectively uses Transformer-based models for visual tracking. wang2024deep explores multi-modal fusion for multi-object tracking. Different from the above, hybrid tracking methods perform an information fusion on various data, such as depth datajiang2019hierarchical , infrared datazhang2020object ; tang2023exploring /multi-modal infrared datacai2023multi , and hyperspectral datachen2024sense , for robust visual object tracking.
@@
Compared with the conventional tracking methods, event-based tracking methods barranco2018real ; camunas2017event ; glover2017robust , which fuse the event data, show their superiority under fast motions and HDR scenes. The recently proposed event-based tracking methods have achieved state-of-the-art performance by leveraging asynchronous tracking-by-detection chen2019asynchronous , deep regression networks chen2020end , spiking Transformers zhang2022spiking , graph embeddingzhu2022learning , bio-inspired unsupervised learningzheng2023spike , data-driven feature trackingmessikommer2023data , and high-rank augmentationzhu2023cross . Other notable event-based computer vision tasks include motion estimation zuo2024cross , visual odometry hidalgo2022event , detection li2022asynchronous , and motion segmentation chen2024segment . From these event-based studies, we can see that the bio-inspired event camera can provide an effective and efficient way to solve various challenging motion-related vision problems, by exploiting and fusing the event data.
@@
However, in the current event-based studies, most methods usually handle the fundamental event-based data association problem in implicit ways, which are designed for their specific tasks. As a result, event-based data association has not been effectively solved by the current event-based works. There are relatively few works focusing on this problem. As one of the pioneer works in this field, zhu2017event proposes a probabilistic data association method based on the event-based features. Also, gallego2018unifying ; gallego2019focus ; peng2022globally present contrast maximization frameworks to implicitly perform data association. The retinal events are usually triggered by multiple motions, and they are greatly affected by a large amount of sensor noise. Due to these difficulties, the aforementioned methods use either indirect or implicit strategies to handle the event-based data association problem. In this paper, we propose a unifying event data association approach (EDA), by which we explicitly handle the event-based data association problem through a robust multi-structural model fitting.
@@
Robust model fitting methods can accurately estimate the model instances (e.g., lines, homograph matrices, or fundamental matrices) in data contaminated with a large number of noise and outliers, and they have been applied in various areas. After the model fitting process, the motion information contained in the event data are naturally and preciously associated and fused for high-level computer vision tasks, such as visual object tracking. The most representative work in robust model fitting is the well-known RANSAC algorithm fischler1981random . To improve the fitting performance, many robust fitting methods that use energy functions pham2014interacting ; barath2018graph , or use preference analysis xiao2020deterministic ; lin2022co , or use hypergraphs xiao2016hypergraph ; wang2019searching ; lin2019hypergraph , or use quantum theorydoan2022hybrid ; farina2023quantum , or use geometric priors xiao2019superpixel ; ma2019locality and clues lin2024robust ; xiao2024latent , have been developed. However, there is no robust model fitting method proposed for event data. In this paper, we are the first to introduce robust model fitting to explicitly target the event association problem for tracking.
3 Event Data Association
@@ 3.1 Overview
In this section, we describe all the components of the proposed EDA approach. The pipeline of the proposed EDA is illustrated in Fig. 2. First, we describe the sequential retinal events, and introduce an asynchronous fusion phase to gather the sequential retinal events, as illustrated in Fig. 2(a)-(d). Then, we present the model hypothesis generation process, to effectively generate model hypotheses, as illustrated in Fig. 2(e)-(f). Finally, we give the details of the model hypothesis selection, to robustly perform the final data association, as illustrated in Fig. 2(g)-(h).
@@ 3.2 Grouping Retinal Events
In this work, our goal is to calculate the correct event trajectories to associate the emitted retinal events. Since these event trajectories are approximately 3D lines over short time intervals (or small retinal event displacements), as pointed out by gallego2018unifying , we need to determine the short time intervals over which we fuse the sequential retinal events $\mathcal{E}$. Currently, event-based methods can be coarsely categorized into two types, i.e., event-by-event-based and groups-of-events-based, as discussed in gallego2022event . A single event only contains very limited information, thus event-by-event-based methods usually rely on additional appearance information. In contrast, many event-based methods use groups of events as their inputs. However, these groups-of-events-based methods usually set the time interval to a constant value (or use a fixed event number as the threshold), which is not adaptive to different scenes. Moreover, fusing events over a constant time fails to leverage the asynchronous nature of the event data.
@@
To keep the asynchronous nature of event data for benefiting high-level event-based tasks, we propose to leverage the asynchronous event-to-frame strategy in chen2019asynchronous to asynchronously fuse the sequential retinal events $\mathcal{E}$ into different sets for data association. According to that strategy, the sequential retinal events are accumulated to form an Adaptive Time-Surface with Linear Time Decay (ATSLTD) frame $\mathcal{F}$. Each of the incoming events will trigger an update at the corresponding pixel coordinates of the ATSLTD frame $\mathcal{F}$, thereby increasing the amount of information contained in the ATSLTD frame $\mathcal{F}$. Then, the amount of information in the ATSLTD frame is measured by the image entropy of $\mathcal{F}$. Here, the image entropy of $\mathcal{F}$ is quantitatively calculated by a Non-Zero Grid Entropy (NZGE) metric. After that, a set of image entropy values is collected from those ATSLTD frames that contain clear and sharp object contours. Here the clear and sharp object contours indicate relatively small event displacements, which keep the event trajectories straight. Finally, a confidence interval $[\alpha,\beta]$ of the collected image entropy values is calculated by using the Student’s t-distribution. During the recording, the emitted sequential retinal events $\mathcal{E}$ will continuously update the ATSLTD frame $\mathcal{F}$. If the image entropy value of the ATSLTD frame $\mathcal{F}$ falls into the interval $[\alpha,\beta]$, the historical retinal events are accumulated for further processing.
@@
Suppose that the image entropy value of the accumulated retinal events collected from time $t_{k}$ falls into the calculated confidence interval $[\alpha,\beta]$ at time $t_{k}+T$. The corresponding accumulated retinal events are $\mathcal{E}_{t_{k},t_{k}+T}$, and the next time interval starts from time $t_{k}+T$. By repeating this process, we can asynchronously fuse the incoming retinal events into different sets for continuous event trajectory estimation and event-based data association.
@@ 3.4 Selecting Model Instances
Then, based on the residual ${\mathcal{R}_{{t_{k}},{t_{k}}+T}}$, for each of the model hypotheses in ^${\hat{\mathcal{L}}_{{t_{k}},{t_{k}}+T}}$, the accumulated retinal events $\mathcal{E}_{t_{k},t_{k}+T}$ can be divided into two sets: i.e., inliers $\mathcal{I}_{t_{k},t_{k}+T}$ and outliers $\mathcal{O}_{t_{k},t_{k}+T}$. For a reference model hypothesis in ^${\hat{\mathcal{L}}_{{t_{k}},{t_{k}}+T}}$, if the residual (i.e., the 3D voxel-to-line distance) for a retinal event is less than an inlier noise scale $\tau$ (which can be estimated by using a robust scale estimator, such as IKOSE wang2011simultaneously ), the corresponding retinal event is considered as an inlier. Otherwise, the corresponding retinal event is considered as an outlier. If the inlier number of the reference model hypothesis is no more than two (i.e., only including the start voxel and the end voxel), this model hypothesis will be considered as noise and removed. After that, the reference model hypothesis will be weighted by measuring the dispersion of its inliers $\mathcal{I}_{t_{k},t_{k}+T}$ on the time axis, which can be written as follows:
@@ 4.1 Parameter Settings
For the parameters used in the proposed EDA, we adopt the same confidence interval $[\alpha,\beta]$ as that of ETD chen2019asynchronous for fair competition. The number of the time slices $N_{s}$ is empirically set to 10 to achieve the trade-off between accuracy and efficiency. The inlier noise scale $\tau$ is fixed to 0.01 for all the experiments.

### v3

# event camera; data association; object tracking; model fitting

@@ Abstract
Event-based approaches, which are based on bio-inspired asynchronous event cameras, have achieved promising performance on various computer vision tasks. However, the study of the fundamental event data association problem is still in its infancy. In this paper, we propose a novel Event Data Association approach (called EDA) to explicitly address the data association problem. The proposed EDA seeks for event trajectories that best fit the event data, in order to perform unifying data association. In EDA, we first asynchronously gather the event data, based on its information entropy. Then, we introduce a deterministic model hypothesis generation strategy, which effectively generates model hypotheses from the gathered events, to represent the corresponding event trajectories. After that, we present a two-stage weighting algorithm, which robustly weighs and selects true models from the generated model hypotheses, through multi-structural geometric model fitting. Meanwhile, we also propose an adaptive model selection strategy to automatically determine the number of the true models. Finally, we use the selected true models to associate the event data, without being affected by sensor noise and irrelevant structures. We evaluate the performance of the proposed EDA on the remote object tracking task. The experimental results show the effectiveness of EDA under challenging scenarios, such as high speed, motion blur, and high dynamic range conditions.
@@
event camera; data association; object tracking; model fitting
[…]
@@ 1 Introduction
As one of the most delicate neural systems, biological retinas can precisely and efficiently capture object motions Ölveczky et al. (2003). Inspired by the biological retina, asynchronous event cameras, such as DVS Lichtsteiner et al. (2008), DAVIS Brandli et al. (2014) and ATIS Posch et al. (2011), have been proposed to mimic it. These bio-inspired silicon retina vision sensors have a very high dynamic range (higher than 100 dB) and an ultra-low latency (less than 1 ms). In particular, each pixel of the event camera can independently and asynchronously emit binary (i.e., On or Off) retinal events in responding to pixel intensity changes in the current environment. Pixel intensity changes are usually triggered by object and camera motions, under relatively smooth external illumination variations. Thus, the retinal events contain clear motion information. Moreover, thanks to the high dynamic range and low latency, event cameras are not affected by those challenges that inherently affect conventional frame-based cameras, such as motion blur and pixel saturation. As a result, event cameras are well suited for motion-related computer vision tasks, such as remote visual tracking.
@@
Event-based methods have achieved promising performance on various tasks Gallego et al. (2020). However, the study of the fundamental event data association problem is still challenging and in its infancy. Unlike a traditional camera, an event camera only sparsely emits binary (i.e., On and Off) retinal events at the edges within a scene, in responding to camera and object motions. Therefore, without the pixel intensity information, it is difficult to calculate event correspondences (i.e., event trajectories) in the temporal domain for event-based data association. Since data association is the key to various fundamental tasks (such as object tracking), the problem of event-based data association needs further investigation.
As the pioneering work on event-based data association, Gallego et al. (2018) proposes a contrast maximization framework to implicitly handle the data association problem based on the corresponding motion parameters. In that framework, the initial motion parameters for each event are empirically set to a range of predefined values for optimization. Since the predefined motion range may not cover all possible event motions, the final result may not be the optimal estimate.
However, Gallego et al. (2018) and other event-based data association methods Zhu et al. (2017); Gallego et al. (2019) show that the events triggered by the same edge in the scene can be associated with each other using an event trajectory. Furthermore, they also show that the event trajectories, triggered at different edges of the scene, are nearly 3D lines over a short time interval in the spatio-temporal domain (see an example in Fig. 1). Based on these observed facts, we explicitly formulate the event-based data association problem as a 3D event trajectory estimation problem in the spatio-temporal domain. Then, a novel event-based data association approach (called EDA) is proposed to perform explicit and accurate event-based data association. In the proposed EDA, the event trajectory estimation problem is effectively solved through the theory of robust multi-structural model fitting. To the best of our knowledge, it is the first time to introduce robust model fitting for event-based data association. The estimated multiple structures correspond to the event trajectories triggered by different motions, and they are used to associate the event data. To evaluate the effectiveness of the proposed EDA, we use EDA to deal with the problem of event-based data association in remote object tracking scenarios. Overall, this paper makes the following contributions:
  • We introduce a deterministic model hypothesis generation strategy to effectively generate model hypotheses in the spatio-temporal domain, spanned by retinal events. The generated model hypotheses are considered as event trajectory candidates for event data association.
@@ 2 Related Work
As mentioned previously, event cameras have shown their great potential on various computer vision tasks Gallego et al. (2020). Different from conventional cameras that are synchronized to the global camera shutter, these bio-inspired vision sensors work in an asynchronous way with ultra low microsecond latencies in responding to object motions, which is inspired by the biological retina.
@@
Thanks to the asynchronous nature, the high spatio-temporal resolution, and the superior HDR properties of event cameras, various event-based methods have been proposed for many computer vision tasks, especially for object tracking. In the field of visual remote tracking, conventional methods can be roughly categorized into three groups: UAV tracking methodsXu et al. (2024); Yang et al. (2024); Huang et al. (2024); Jiang and Yin (2023), RGB tracking methodsYe et al. (2024); Liang et al. (2024); Wang et al. (2023); Hou et al. (2023); Liang et al. (2021), hyperspectral tracking methodsGao et al. (2025); Guo et al. (2024), and RGB-T tracking methodsXue et al. (2023). Compared with conventional methods, event-based tracking methods Barranco et al. (2018); Camuñas-Mesa et al. (2017); Glover and Bartolozzi (2017) show their superiority under fast motions and HDR scenes. The recently proposed event-based tracking methods have achieved state-of-the-art performance by leveraging motion compensationMitrokhin et al. (2018), asynchronous tracking-by-detection Chen et al. (2019), and deep regression networks Chen et al. (2020). Other notable motion-related event-based tasks include optical flow estimation Zhu et al. (2019), camera motion estimation Gallego et al. (2017), and motion segmentation Stoffregen et al. (2019). In addition, there are also some works in relation to event-based feature extraction Clady et al. (2015), classification Lagorce et al. (2017), and stereo vision Stoffregen et al. (2019). From these event-based studies, we can see that the bio-inspired event camera can provide an effective and efficient way to solve various challenging vision problems, by exploiting the event data.
@@
However, in the current event-based studies, most methods usually handle the fundamental event-based data association problem in implicit ways, which are designed for their specific tasks. As a result, event-based data association has not been effectively solved by the current event-based works. There are relatively few works focusing on this problem. As one of the pioneer works in this field, Zhu et al. (2017) proposes a probabilistic data association method based on the event-based features. Also, Gallego et al. (2018, 2019) present a unifying contrast maximization framework to implicitly perform data association. The retinal events are usually triggered by multiple motions, and they are greatly affected by a large amount of sensor noise. Due to these difficulties, the aforementioned methods use either indirect or implicit strategies to handle the event-based data association problem. In this paper, we propose a unifying event data association approach (EDA), by which we explicitly handle the event-based data association problem through a robust multi-structural model fitting.
@@
Robust model fitting methods can accurately estimate the model instances (e.g., lines, homograph matrices, or fundamental matrices) in data contaminated with a large number of noise and outliers, and they have been applied in various areas. The most representative work in robust model fitting is the well-known RANSAC algorithm Fischler and Bolles (1981). To improve the fitting performance, many robust fitting methods that use energy functions Pham et al. (2014); Barath and Matas (2018), or use preference analysis Magri and Fusiello (2014, 2019); Xiao et al. (2020), or use hypergraphs Xiao et al. (2016); Wang et al. (2019), or use geometric priors Xiao et al. (2019); Ma et al. (2019, 2021), have been developed. However, there is no robust model fitting method proposed for event data. In this paper, we are the first to introduce robust model fitting to explicitly target the event association problem for object tracking.
3 Proposed EDA Method
@@ 3.1 Overview
In this section, we describe all the components of the proposed EDA approach. The pipeline of the proposed EDA is illustrated in Fig. 2. First, we describe the sequential retinal events, and introduce an asynchronous gathering phase to gather the sequential retinal events, as illustrated in Fig. 2(a)-(d). Then, we present the model hypothesis generation process, to effectively generate model hypotheses, as illustrated in Fig. 2(e)-(f). Finally, we give the details of the model hypothesis selection, to robustly perform the final data association, as illustrated in Fig. 2(g)-(h).
@@ 3.2 Grouping Retinal Events
In this work, our goal is to calculate the correct event trajectories to associate the emitted retinal events. Since these event trajectories are approximately 3D lines over short time intervals (or small retinal event displacements), as pointed out by Gallego et al. (2018), we need to determine the short time intervals over which we gather the sequential retinal events $\mathcal{E}$. Currently, event-based methods can be coarsely categorized into two types, i.e., event-by-event-based and groups-of-events-based, as discussed in Gallego et al. (2020). A single event only contains very limited information, thus event-by-event-based methods usually rely on additional appearance information. In contrast, many event-based methods use groups of events as their inputs. However, these groups-of-events-based methods usually set the time interval to a constant value (or use a fixed event number as the threshold), which is not adaptive to different scenes. Moreover, gathering events over a constant time fails to leverage the asynchronous nature of the event data.
@@
To keep the asynchronous nature of event data for benefiting high-level event-based tasks, we propose to leverage the asynchronous event-to-frame strategy in Chen et al. (2019) to asynchronously gather the sequential retinal events $\mathcal{E}$ into different sets for data association. According to that strategy, the sequential retinal events are accumulated to form an Adaptive Time-Surface with Linear Time Decay (ATSLTD) frame $\mathcal{F}$. Each of the incoming events will trigger an update at the corresponding pixel coordinates of the ATSLTD frame $\mathcal{F}$, thereby increasing the amount of information contained in the ATSLTD frame $\mathcal{F}$. Then, the amount of information in the ATSLTD frame is measured by the image entropy of $\mathcal{F}$. Here, the image entropy of $\mathcal{F}$ is quantitatively calculated by a Non-Zero Grid Entropy (NZGE) metric. After that, a set of image entropy values is collected from those ATSLTD frames that contain clear and sharp object contours. Here the clear and sharp object contours indicate relatively small event displacements, which keep the event trajectories straight. Finally, a confidence interval $[\alpha,\beta]$ of the collected image entropy values is calculated by using the Student’s t-distribution. During the recording, the emitted sequential retinal events $\mathcal{E}$ will continuously update the ATSLTD frame $\mathcal{F}$. If the image entropy value of the ATSLTD frame $\mathcal{F}$ falls into the interval $[\alpha,\beta]$, the historical retinal events are accumulated for further processing.
@@
Suppose that the image entropy value of the accumulated retinal events collected from time $t_{k}$ falls into the calculated confidence interval $[\alpha,\beta]$ at time $t_{k}+T$. The corresponding accumulated retinal events are $\mathcal{E}_{t_{k},t_{k}+T}$, and the next time interval starts from time $t_{k}+T$. By repeating this process, we can asynchronously gather the incoming retinal events into different sets for continuous event trajectory estimation and event-based data association.
@@ 3.4 Selecting Model Instances
Then, based on the residual ${\mathcal{R}_{{t_{k}},{t_{k}}+T}}$, for each of the model hypotheses in ^${\hat{\mathcal{L}}_{{t_{k}},{t_{k}}+T}}$, the accumulated retinal events $\mathcal{E}_{t_{k},t_{k}+T}$ can be divided into two sets: i.e., inliers $\mathcal{I}_{t_{k},t_{k}+T}$ and outliers $\mathcal{O}_{t_{k},t_{k}+T}$. For a reference model hypothesis in ^${\hat{\mathcal{L}}_{{t_{k}},{t_{k}}+T}}$, if the residual (i.e., the 3D voxel-to-line distance) for a retinal event is less than an inlier noise scale $\tau$ (which can be estimated by using a robust scale estimator, such as IKOSE Wang et al. (2011)), the corresponding retinal event is considered as an inlier. Otherwise, the corresponding retinal event is considered as an outlier. If the inlier number of the reference model hypothesis is no more than two (i.e., only including the start voxel and the end voxel), this model hypothesis will be considered as noise and removed. After that, the reference model hypothesis will be weighted by measuring the dispersion of its inliers $\mathcal{I}_{t_{k},t_{k}+T}$ on the time axis, which can be written as follows:
@@ 4.1 Parameter Settings
For the parameters used in the proposed EDA, we adopt the same confidence interval $[\alpha,\beta]$ as that of ETD Chen et al. (2019) for fair competition. The number of the time slices $N_{s}$ is empirically set to 10 to achieve the trade-off between accuracy and efficiency. The inlier noise scale $\tau$ is fixed to 0.01 for all the experiments.